An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 798 hours and a standard deviation of 80 hours. Find the probability that a random samples of 39 bulbs will have an average life of less than 814 hours


1
Expert's answer
2022-02-02T16:20:49-0500

"P(X<814)=P(Z< \frac{814-798}{ \frac{80}{ \sqrt{39}

}})"


"P(X<814)=P(Z<1.25)"


From using normal distribution table, probability is:


"P(X<814)=0.8944"


There is 89.44% probability that a random samples of 39 bulbs will have an average life of less than 814 hours.